# VideoHallu: Evaluating and Mitigating Multi-modal Hallucinations for Synthetic Videos

### v1

@@ 4 Experiment and Results
In this section, we evaluate 16 SoTA MLLMs on our curated test dataset (Table 1). For models not trained with reinforcement learning, we use standard prompting to elicit direct answers. For those trained with reinforcement learning or chain-of-thought supervised finetuning (e.g., Video-R1-cot [9] and VideoChat-R1-think [10]), we prompt them to reason step by step by analyzing key video frames before generating a final answer. Figure 4 highlights hallucinations produced by SoTA models across all four categories in synthetic video understanding tasks, with the hallucinated contexts marked within each answer. Additional examples are included in Appendix A.
Evaluation:
LLM-as-a-judge [66, 67, 68] has shown promising improvement and high correlation with human judgments compared to previous work for simple and easy gold and generated answer comparison. Since our answers include free-form and open-eneded question-answer pairs, we use GPT-4-as-a-Judge to judge the correctness of model generated response with our written gold answers. We extract the final answer from Video-R1-cot and Videochat-R1-think as the generated response for GPT-4 to judge. To ensure the robustness and correctness of LLM-as-a-Judge in our dataset, we randomly sample 200 answer pairs with length greater than five words and annotate their correctness, with a 99.3% human-GPT agreements.
@@ 4.1 Results and Analysis
MLLMs struggle to detect counterintuitive phenomena and abnormalities in generated videos.
Qwen-VL series (7B/32B) achieves the highest overall accuracy at around 51% (Table 1). However, even the best models often fail to answer simple visual questions about synthetic videos, relying heavily on language priors or prior knowledge from real-world videos, which leads to unfaithful or incorrect assumptions from real-life data (Figure 5). Qwen2.5-VL’s advantage stems from broader training dataset coverage, a vision encoder tailored for video tasks, and strong cross-modal alignment. In contrast, models like GPT-4o treat video inputs as frame sets without modeling temporal dynamics, falling behind other competitive SoTA models–Gemini-2.5-pro and Gemini-2.5-Flash.
Despite Qwen2.5-VL’s lead, the overall accuracy remains below 60% across all models. The performance across sub-categories is uneven: models excel at tasks requiring direct observation (e.g., instruction-following, alignment, spatial-temporal consistency), but struggle with reasoning-intensive questions involving common sense or physics knowledge. Within each category, perception-oriented tasks like entity recognition yield higher accuracy than reasoning-heavy ones like spatial relations or entity counting, where language priors and visual priors from real-life videos often introduce hallucinations. This highlights current models’ limited capacity for integrating memorized knowledge into multi-modal reasoning.
Larger models tend to achieve higher accuracy, but their performance is ultimately constrained by the capabilities of the vision encoder.
We evaluate models across a range of sizes, from small open-source models like InternVL3-2B to large closed-source ones like Gemini-2.5-Pro and GPT-4o. Generally, larger models demonstrate stronger language and visual reasoning abilities on all categories Our results confirm this trend: within the same series, such as Qwen and InternVL series, larger models outperform smaller ones across tasks. However, for models within the same series (Qwen2.5-VL, InterVL3), accuracy gains reach peak on synthetic videos once the model size exceeds 7B, showing that further improvements in video understanding rely less on scaling the language model and more on enhancing the vision encoder and the quality of training data. In addition, overall performance is also highly relevant to how models perceive video inputs. GPT-4o tends to treat videos frame-by-frame, aiding frame-level alignment checks but harming inter-frame motion tracking, which again shows the importance of vision encoder architecture in video understanding.
@@
Figure 4: Hallucination showcases for SoTA models on VideoHallu. We collect hallucination cases observed during SoTA MLLM evaluations on synthetic video tasks. Each example includes the generation prompt, key frames, questions, human-annotated ground truth, and hallucinated answers from GPT-4o, Qwen2.5-VL, and Gemini-2.5-Pro, with hallucinations marked in Red. Video examples are available (here).
@@
4.2 Does Chain-of-Thought Reinforcement Learning Help Synthetic Video Understanding?
Chain-of-thought reasoning encourages models to rely more on prior language knowledge, often at the expense of visual grounding, which increases the risk of hallucinations.
Despite the recent success of the DeepSeek series [13], where Reinforcement Fine-Tuning (RFT) methods like GRPO [71] have enhanced reasoning in small models like Qwen2-VL-7B on math problems, our results show limitations of chain-of-though RL algorithms. In Table 1, we evaluate two RFT-trained video understanding models, Video-R1 [9] and VideoChat-R1 [10], under different prompt styles: short-answer to directly generate the answers (Video-R1, VideoChat-R1) and chain-of-thought to first think step by step then generate the final answer (Video-R1-CoT, VideoChat-R1-think). Both RFT models underperform their base model (Qwen2.5-VL-7B), and short-answer prompt style outperform those explicitly prompted for reasoning. This suggests that current RFT approaches, often based on limited, homogeneous data, boost performance on in-distribution tasks (e.g., real-world video understanding) but fail to generalize to synthetic videos. While chain-of-thought RFT can enhance performance on reasoning-heavy tasks like math problems, it is less effective for synthetic video understanding, where videos often depict counterintuitive or abnormal phenomena that conflict with real-life phenomenons. In such cases, encouraging extensive reasoning can cause the LLM backbone to rely too heavily on real-world commonsense knowledge, overlooking synthetic visual evidence and leading to hallucinated responses. GRPO-based RFT further amplifies this issue by encouraging models to develop their own reasoning patterns using weak supervision and non-semantic rewards like ROUGE [72], which poorly correlate with human judgment [73, 68]. This misalignment creates a "perception gap" between model and human reasoning, often resulting in hallucinations—particularly on reasoning-heavy questions—explaining the weaker performance of chain-of-thought RFT models on our benchmark.
4.3 Can MLLMs learn counter-intuitive commonsense knowledge from synthetic data and curriculum learning?
Although Video-R1 is trained on massive number of diverse datasets, it still falters on synthetic video understanding, or even show lower to no improvement on VideoHallu. In this section, we use curriculum learning [33] to train Qwen2.5-VL-7B on three datasets in the order of difficulty– 4,500 from the video QA subsection of Video-LLaVA, 1,000 from PhysBench [22], and 800 from VideoHallu. We aim to study two research questions: 1. Can we improve MLLMs’ reasoning and understanding abilities on synthetic videos through reinforcement learning finetuning? 2. Does curriculum learning help video understanding? We combine the training dataset for GRPO training from the following three sources: One partition of VideoHallu dataset, the subsection of the open-end question-answering dataset of Video-LLaMA, along with the videos with physical law knowledge from PhysBench [22]. The combination of the fine-tuning dataset is inspired by the fine-tuning procedure of Video-R1 [9], where we try to improve the models’ video understanding abilities over common sense and physical laws by designing the entire training process in a curriculum manner with three datasets.
General-Real-World (GRW):
The dataset for real-world video question-answering (Video-LLaMA [34]), to improve the base models’ ability in video understanding. We sample 3,000 videos from this dataset as the training set.
Physics-Real-World (PRW):
PhysBench [22] focuses on real-world common and physical knowledge, while this stage is to assist the model to understanding the indications and procedures, further refine the representations of physical and common sense-related knowledge, with the language prior within the video understanding model. We sample 1,000 videos from the dataset as the training set.
Synthetic Reasoning (SR):
The dataset contains all synthetic videos question-answering (VideoHallu), focusing on common and physical knowledge. This stage helps the model overcome potential hallucinations within the language priors when reasoning over synthetic video and become more sensitive in probing common sense and physics violations in synthetic videos. Our training set size is 800 video QA pairs.
Specifically, we use GRPO to train Qwen2.5-VL-7B on three datasets separately first– GRW-R1-7B, PRW-R1-7B, SR-R1-7B. We use Answer-Equivalence BERT [68], which has been fine-tuned specifically on answer-correctness pairs, to evaluate the correctness of generated answers compared to the gold answer. The reward is defined as the similarity score between the embedding representations of the predicted answer $a_{pred}$ and the gold answer $a_{gold}$, calculated similarly to BERTScore [74] as follows:
@@ 4.3 Can MLLMs learn counter-intuitive commonsense knowledge from synthetic data and curriculum learning?
, where $E(a_{gold})$ and $E(a_{pred})$ denote the embedding vectors of the predicted and gold answers from the finetuned BERT, respectively. We train each model group with one epoch with gradient accumulation step of 1. We use the saved checkpoints for GRW and PRW and then train the model on the synthetic dataset to perform curriculum learning. Our results (Table 2) show that only training on general-world data does not improve model’s understanding abilities on synthetic video understanding, but adding physics knowledge can improve synthetic video understanding robustness mainly in the physics domain. In addition, we show that curriculum learning to first learn general physics knowledge from the data, then from synthetic videos can help models gain most understanding on synthetic video understanding.
@@ A-EC A-EP A-ERAC A-SR SC-CD SC-SD SC-TD CS-K CS-R P-C P-CAP P-M P-ST
Figure 5: Evaluation Breakdown of Fine-Tuned Models. We show results for (a) previous SoTA MLLMs, (b) models fine-tuned on sub-datasets, and (c) models fine-tuned on the full dataset via curriculum learning. Compared to the baseline (Qwen2.5-VL-7B), reinforcement fine-tuning on commonsense and physics data improves models’ reasoning and overall performance in synthetic video understanding.
@@ 4.4 Discussions
Throughout the entire evaluations over our benchmark and the reinforcement fine-tuning over pre-trained MLLMs, we gather essential insights to accelerate further improvement over future MLLMs for synthetic video understanding. We list them as follows:
1. Visual encoders constrain MLLMs’ performance in synthetic video understanding.
In Section 4.1, all MLLMs evaluated suffer from hallucination in synthetic video tasks, achieving poor overall accuracy below 52% and consistently performing better on perception-oriented tasks (e.g., alignment, spatial-temporal consistency) than on reasoning-heavy tasks involving commonsense and physical reasoning. This gap partly stems from how MLLMs perceive videos: current visual encoders struggle to capture context-level temporal and spatial dynamics. Instead of flexibly encoding entity-level, semantic-sensitive information critical for synthetic video understanding, MLLMs tokenize videos into patches or fixed-length clips, inherited from MLLMs’ visual encoders, which limits their ability to handle the non-realistic, dynamic patterns common in synthetic videos.
2. Hallucination in synthetic video understanding stems from language priors, which makes detecting synthetic visual abnormalities worse.
We observe severe hallucinations in the reasoning processes of synthetic video understanding. SoTA models struggle at reasoning-oriented QA tasks, and RFT only on real-world data even worsens performance (Table 1). This stems from current MLLMs’ strong language priors, learned from datasets dominated by real-world videos with explicitly represented contexts, where commonsense and physical laws are typically implicit. Despite seeing annotated synthetic examples during training, models assume these principles as default truths without truly learning from visual cues. As a result, when faced with generated videos that contain content conflicting with real-world phenomena, models struggle to reconcile the visual input with their prior knowledge. Furthermore, reasoning built on these flawed understandings leads to even more severe hallucinations, as shown in Video-R1-cot and VideoChat-R1-think (Table 1). We propose that future MLLMs for synthetic or physical-reasoning video tasks should focus on refining reasoning abilities to better leverage language priors while actively correcting hallucinated vidual understandings.
@@
In Section 4.2, we show that RFT-enhanced models like Video-R1 and VideoChat-R1 do not outperform their base models (Qwen2.5-VL-7B) on synthetic video understanding tasks. However, in Section 4.3, we demonstrate that reinforcement fine-tuning methods like GRPO, when combined with a curated dataset covering both general and physics-specific video understanding across real and synthetic videos can improve model performance. Our results reveal that it is the quality and coverage of the data—not just the fine-tuning method—that drive improvements. With both positive and negative examples, clearly annotated reasoning procedures, and reasoning-stimulating post-training like GRPO, even small-scale models like Qwen2.5-VL-7B achieve QA accuracy gains on validation sets. We show the importance of combining high-quality data and reasoning-focused fine-tuning to enhance synthetic video understanding in future models, encouraging models to fundamentally grasp commonsense and physics knowledge and apply them in reasoning about more advanced concepts.
@@ 5 Conclusion
We introduce VideoHallu, a novel benchmark targeting hallucination probing and mitigation in synthetic video understanding. We curate expert-annotated, diverse, reasoning-heavy QA pairs covering alignment, spatial-temporal consistency, commonsense, and physics reasoning to probe hallucinations induced by real-world training priors in MLLMs that are not generalized to synthetic generated videos. Evaluation on SoTA MLLMs show hallucination and poor performance on synthetic videos. We fine-tune MLLMs using GRPO with real-world video data, physics reasoning tasks, and VideoHallu synthetic videos organized via curriculum learning, achieving accuracy improvements. Our results show the importance of incorporating physics and commonsense reasoning data when refining MLLMs for synthetic video tasks. However, scalability remains a limitation, as producing high-quality annotations is costly. Future work will focus on expanding the dataset from synthetic videos and enhancing reinforcement fine-tuning to foster stronger and more robust reasoning in MLLMs, ultimately advancing solid physical-world understanding in video tasks.
Appendix A Hallucination Showcases

### v2

@@ 3 Experiment and Results
Given the collected adversarial QA pairs, we evaluate 20 SOTA MLLMs (Table 2). For models not trained with reinforcement learning, we use standard prompting to elicit direct answers. For those trained with reinforcement learning or chain-of-thought (CoT) supervised finetuning (e.g., Video-R1-CoT [7] and VideoChat-R1-think [8]), we prompt them to perform step-by-step critical thinking and reasoning before generating a final answer (Appendix. B). Figure 4 highlights hallucinations produced by SoTA models across all four categories in synthetic video understanding tasks, with the hallucinated contexts marked within each answer. Additional examples are included in Appendix D.
Automatic Evaluation:
We adopt LLM-as-a-Judge [30, 31, 32] as our automatic evaluation method, given its strong correlation with human judgments on open-ended QA tasks. GPT-4 is used to assess the correctness of model responses, using final answers from Video-R1-CoT and VideoChat-R1-think. To validate reliability on our dataset, we manually annotated 200 randomly sampled answer pairs, achieving 99% agreement with GPT-4.
@@ 3.1 Results and Analysis
MLLMs struggle with counterintuitive phenomena and abnormalities in generated videos.
Despite strong performance on real-world video tasks, SOTA MLLMs achieve less than 52% accuracy on our benchmark (Table 2, Figure 3), performing poorly in commonsense and physical reasoning. Models like Qwen2.5-VL (7B/32B), GPT-4o, and Gemini-2.5-Pro fail to detect abnormalities in synthetic videos and often hallucinate due to their language priors. As shown in Figure 4, SOTA MLLMs consistently fail to answer counterintuitive questions involving physical abnormalities, as none of them correctly identify the implausibility of a shattered watermelon reassembling. They also overlook abrupt, counterfactual entity changes, such as a quail suddenly turning into a rooster in the spatial-temporal consistency case. These errors highlight their inability to apply critical thinking for abnormality reasoning based on actual visual evidence rather than textual priors.
@@
Figure 4: Hallucination showcases for SOTA models on VideoHallu. We collect hallucination cases observed during SOTA MLLM evaluations on synthetic video tasks. Each example includes the generation prompt, key frames, questions, human-annotated ground truth, and hallucinated answers from GPT-4o, Qwen2.5-VL, and Gemini-2.5-Pro, with hallucinations marked in Red.
@@
Chain-of-thought learned from real-world videos does not help synthetic video understanding.
MLLMs enhanced by Reinforcement Fine-Tuning (RFT), such as GRPO [40] used in the DeepSeek series [11], show a great potential in improving reasoning and critical thinking abilities over small base models like Qwen2.5-VL-7B in some reasoning-heavy tasks like mathematics. This provides an alternative method for commonsense/physics abnormality detection in synthetic videos. In Table 2, we evaluate two RFT-trained models, Video-R1 [7] and VideoChat-R1 [8], under the short-answer prompt style that directly generates the answers (Video-R1, VideoChat-R1), and chain-of-thought prompt style where MLLMs first think step by step then generate the final answer (Video-R1-CoT, VideoChat-R1-think). Surprisingly, both RFT models underperform their base model (Qwen2.5-VL-7B), and short-answer prompts outperform CoT prompts, showing little improvement on counterintuitive and critical thinking questions (Figure 3). This indicates that current RFT methods, often trained on real-world video data, generalize poorly to synthetic videos.
Solely pre-training on real-world data biases visual grounding.
While RFT improves reasoning on math and real-world videos, it struggles with counterintuitive synthetic content that contradicts real-world norms and lacks video-grounded critical thinking. In such cases, chain-of-thought prompting can bias the LLM backbone to rely too heavily on prior commonsense knowledge, neglecting synthetic visual cues and leading to hallucinated responses. For instance, in the third case in Figure 4, VideoChat-R1-think responds: “The video shows a feather and a rock being dropped… This is a classic demonstration of Galileo’s principle…”, a plausible explanation grounded in language priors, but misaligned with the actual video. The model generates incorrect conclusions and hallucinated reasoning, illustrating how chain-of-thought prompting may amplify reliance on priors and increase hallucination risk in synthetic video understanding.
3.2 Learning Reasoning from Synthetic Data
Section 3.1 shows that current MLLMs struggle with counterintuitive questions in synthetic videos, often hallucinating and failing to perform critical thinking for abnormality detection on VideoHallu, despite their extensive pre-training datasets. This section investigates the key question: Can MLLMs learn counter-intuitive commonsense knowledge and improve critical thinking in detecting abnormalities in synthetic videos?
A natural approach is to use RFT to enhance SOTA MLLMs. However, a key challenge lies in the gap between standard alignment-based video QA tasks used in pre-training and the abnormality reasoning required for synthetic videos. To address this, we propose a three-stage curriculum training pipeline using: (a) General-Real-World (GRW): 3,000 video QA pairs from Video-LLaMA [27] to boost general video understanding. (b) Physics-Real-World (PRW): 1,000 real-world videos QA pairs from PhysBench [28], targeting physical and commonsense reasoning to refine model representations. (c) Synthetic Reasoning (SR): 800 synthetic video QA pairs from VideoHallu, focusing on commonsense and physics abnormality detection, aiming to reduce language-prior-induced hallucinations and boost abnormality detection.
We use Qwen2.5-VL-7B as the base model and enhance it with GRPO through reinforcement fine-tuning (RFT) across the three dataset partitions in a curriculum learning setup. During fine-tuning, we employ Answer-Equivalence BERT [32], trained on answer-correctness pairs, to assess the accuracy of generated responses. The RFT reward is defined as the similarity between the embeddings of the predicted answer $a_{pred}$ and the gold answer $a_{gold}$, computed in a BERTScore-like manner [41]:
@@ 3.2 Learning Reasoning from Synthetic Data
where $\mathbf{E}(a_{gold})$ and $\mathbf{E}(a_{pred})$ denote the embedding vectors of the predicted and gold answers from the finetuned BERT, respectively.
We present our RFT results under the curriculum learning setup combining GRW/PRW with SR (GRW+SR-R1-7B, PRW+SR-R1-7B), along with ablations using individual partitions (GRW-R1-7B, PRW-R1-7B, SR-R1-7B). All models are fine-tuned for one epoch with a gradient accumulation step of 1. Results (Table 3, Figure 5) show that general-world data alone does not enhance performance on synthetic video understanding. However, incorporating physics knowledge notably improves robustness in commonsense/physics abnormality detection. Even with limited data, RFT models show consistent gains in video QA across categories.
@@ A-EC A-EP A-ERAC A-SR SC-CD SC-SD SC-TD CS-K CS-R P-C P-CAP P-M P-ST
Figure 5: Evaluation Breakdown of Fine-Tuned Models. We show results for (a) previous SOTA MLLMs, (b) models fine-tuned on sub-datasets, and (c) models fine-tuned on the full dataset via curriculum learning. Compared to the baseline (Qwen2.5-VL-7B), RFT on commonsense and physics data improves models’ reasoning and overall performance in synthetic video understanding.
@@ 3.3 Discussions
Throughout the evaluations over our benchmark and the reinforcement fine-tuning over pre-trained MLLMs, we gather essential insights to accelerate further improvement over future MLLMs for synthetic video understanding. We list them as follows:
1. MLLMs hallucinate in counterintuitive question answering.
As shown in Table 2 and Figure 3, all tested SOTA MLLMs, including large models like Qwen2.5-VL (7B/32B), GPT-4o, and Gemini-2.5-Pro, as well as smaller models (<7B), struggle with counterintuitive QA on synthetic videos in VideoHallu. MLLMs often default to their embedded commonsense and physics priors, even when prompted to rely on video content (Figure 4). These hallucinations, caused by misalignment between video context and real-world norms, are rare in real-world QA but prevalent in synthetic settings, particularly for counterfactual reasoning. While MLLMs are exposed to some synthetic data during training, their learning is dominated by real-world distributions where physical laws and commonsense are implicitly upheld. Consequently, they treat such rules as universal, failing to adapt when synthetic visuals contradict them, resulting in hallucinated answers when visual evidence clashes with prior knowledge.
2. Critical thinking biased by language priors deteriorates synthetic visual abnormality detection.
In Section 3.1, we present results on RFT-enhanced MLLMs. While RFT improves critical thinking in real-world video QA, all RFT models we tested, such as Video-R1-CoT and VideoChat-R1-think in Table 2, underperform their base model (Qwen2.5-VL-7B) without notable gains on commonsense/physics-oriented questions. We attribute this to flawed critical thinking patterns in RFT-enhanced MLLMs. Although CoT reasoning helps mitigate hallucinations in real-world settings by unpacking complex reasoning, it fails in synthetic video contexts, where detecting abnormalities requires a deep, grounded understanding of real-world commonsense and physics. RFT models often display a shallow understanding of these concepts, making their CoT reasoning less reliable and more prone to hallucinations in counterintuitive scenarios, while hallucinations that CoT reasoning can inadvertently reinforce, making them harder to correct.
@@
Given the need to improve MLLMs’ performance in synthetic video abnormality detection, as shown in Section 3.1, we conduct RFT experiments over SOTA MLLMs that use a curriculum learning pipeline with a composed dataset spanning both general and physics-specific video understanding across real and synthetic videos. Our results show that, after curriculum learning, MLLMs exhibit notable improvements in critical thinking and their ability to handle counterintuitive scenarios. Our results suggest that it is the quality and coverage of the data, not just the fine-tuning method, that drive gains. With a small but well-annotated dataset containing both positive and negative examples, detailed reasoning steps, and reasoning-oriented training like GRPO, even small models like Qwen2.5-VL-7B show improved QA accuracy. This highlights the importance of high-quality, reasoning-rich data in helping MLLMs internalize and apply commonsense and physics knowledge, even with limited post-training resources.
4 Related Work
Hallucinations in MLLMs.
Hallucinations refer to the persistent challenge of generating outputs that contradict or misrepresent the target texts, images, or videos [42, 43]. It arises from conflicts between the language priors of MLLMs and the actual visual inputs [29], which is more severe in video understanding than in static image understanding due to the complex entanglement of spatial-temporal information across the timeline and the contextual cues associated with entities within frames. A line of prior work, such as VideoHallucer [44], EventHallusion [45], and HAVEN [46], established benchmarks for evaluating model hallucination on both entities and events within videos, while also proposing methods to enhance the video understanding capabilities of MLLMs [47, 48]. However, most prior works on hallucination, particularly in the video domain, rely on real-world factual data, rather than synthetic data generated by generative models. Hallucination in generative video understanding models remains an open and largely unexplored research area.
Evaluation of Video Generation Models.
Video generation models [1, 3, 16, 34, 17, 35] are widely used in content creation, robotics, and training [49], but issues like cross-modal misalignment [50] and common sense violations [28] remain major challenges. Benchmarks [51, 52] and model judges [53] are developed to evaluate spatial-temporal consistency, cross-modal alignment, and visual fidelity [18] to address this. Recent work has shifted toward context-level evaluation, emphasizing dynamic properties, physical law compliance [54, 55, 56], and commonsense reasoning in generation prompts. Some approaches use human-annotated data to fine-tune evaluators that detect physical and common sense violations in motion, layout, and state transitions [20, 57]. Despite progress, current methods often lack interpretability and fine-grained reasoning, particularly regarding video abnormalities and critical thinking.
Video understanding for evaluation.
As a core task in computer vision [58, 59], video understanding is traditionally centered on recognizing entities and events across video timelines [60, 61]. Foundation MLLMs [62, 49] have enabled general-purpose video understanding by leveraging large-scale data and high-capacity architectures. Recent work focuses on unifying video and language through instruction tuning [63] and shared representation spaces [27, 39], or scaling vision-language models to video [6, 25, 5]. Given deployment demands, lightweight expert models have been developed for tasks like video QA [64, 65, 36]. Reinforcement learning has also been used during fine-tuning [7, 8] to promote self-improvement and chain-of-thought reasoning. Despite progress, most work targets real-world video QA, with underexplored challenges in detecting commonsense or physical violations and reducing hallucinations in synthetic video QA [66].
@@ 5 Conclusion
We introduce VideoHallu, a benchmark designed to evaluate counterintuitive, commonsense, and physics reasoning in synthetic video understanding. It features expert-annotated, reasoning-intensive QA pairs spanning alignment, spatial-temporal consistency, commonsense, and physics categories to assess MLLMs’ ability to detect abnormalities and violations of physical laws. Evaluation of SOTA MLLMs on VideoHallu reveals persistent hallucinations and critical thinking failures. Fine-tuning with GRPO on a curriculum-organized dataset of physics-focused real and synthetic videos leads to notable accuracy improvements. These results highlight the value of incorporating structured physics and commonsense reasoning data to improve MLLM performance on synthetic video tasks. However, scalability remains a limitation, as generating high-quality annotations and fine-tuning MLLMs at scale is costly, and limited access to data and compute constrains our further progress. Our future work will focus on expanding synthetic video datasets with abnormality QA pairs to train MLLMs for critical, visually-grounded reasoning. Scaling with adversarial QA pairs can enhance robustness and enable automatic video evaluation via prompt decomposition, reducing reliance on human annotations.
Appendix A Video Understanding and Evaluation Categorization/Motivation
In this section, we provide details on specific categorizations of errors video generation models can make. We draw inspiration from basic video quality evaluation definitions from MVBench [9] and WorldModelBench [20] to first organize the current challenges of video generations and evaluations in four basic categories (Figure 2). Given the probing target of each question-answering pair and the demand for reasoning abilities or prior knowledge of the LLM backbone to solve the question provided, we divide the question-answering pairs for testing MLLM-as-evaluators into four major categories with sub-categories.
The categorization is to go beyond superficial metrics like frame consistency or resolution by enabling rigorous evaluation through the identification of visual abnormalities across predefined categories. To achieve this, we design targeted adversarial questions that expose these anomalies. This allows us to assess whether current SOTA MLLMs can effectively detect and interpret such issues, which is an essential step toward scalable and interpretable video evaluation. We further extend these principles to define our video understanding criteria benchmark.
Alignment checks whether the model accurately identifies basic entity details and ensures the video content fully aligns with the prompt without omissions or discrepancies.
Spatial-Temporal Consistency evaluates whether the model can detect smooth, consistent changes in objects, actions, and viewpoints over time, without abrupt or abnormal transitions in space or time.
Common Sense Reasoning assesses the model’s ability to apply general knowledge and reasoning to detect conflicts between common sense and the visual context, ensuring it interprets the prompt correctly without hallucinating entities or actions.
  • Reasoning (CS-R): Tests the model’s ability to interpret problem cues—including emotional or environmental hints, and solve them through reflection and chain-of-thought.
Physics assesses the model’s ability to detect physical inconsistencies, such as violations of gravity, motion dynamics, or conservation laws, requiring careful reasoning about object properties and movements even if not explicitly stated.
  • Motion (P-M): Evaluates the model’s grasp of motion-related physics (like gravity, linear/circular motion, relative movement, and fluid dynamics), spotting inconsistencies or abrupt changes.
Appendix B Prompt Templates
We provide the prompt templates we use for CoT prompt (Table 4) then generate the final answer (Video-R1-CoT and VideoChat-R1-thinking) and prompt templates for generating answers directly (Table 5).
CoT Prompt Template
System Prompt: A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>
Input: Please think about this question as if you were a human pondering deeply. Engage in an internal dialogue using expressions such as ‘let me think’, ‘wait’, ‘Hmm’, ‘oh, I see’, ‘let’s break it down’, etc, or other natural language thought expressions. It is encouraged to include self-reflection or verification in the reasoning process. Provide your detailed reasoning between the <think> </think> tags, and then give your final answer between the <answer> </answer> tags.
Question: {Question}
Table 4: The prompt template for Video-R1-CoT and VideoChat-R1-thinking to generate answers. This prompt encourages them to first think critically about the video and the question then generate a final answer.
Direct Answer Prompt Template
System Prompt: A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant provide answers within the <answer> </answer> tags: <answer> answer here </answer>
Input: You will be given a video and a question. Please provide an answer to the question based on the video enclosed by <answer> your answer </answer> tags.
Question: {Question}
Answer:
Table 5: Direct answer directly prompts a model to generate the answer without generating additional chain-of-thoughts.
Appendix C Common Sense and Video-dependent Question-Answering
Our benchmark, VideoHallu, is designed to evaluate MLLMs’ abilities to detect abnormalities in synthetic videos—a task often confounded by hallucinations stemming from commonsense or physical knowledge embedded in their language priors. This section breaks down model performance across question types in VideoHallu, including:
  • Common Sense-only Questions: These can be answered using language priors alone, without relying on video input. e.g., What typically happens when a bullet hits a watermelon? (Answer: It explodes into pieces.)
  • Counterintuitive Questions: Target counterfactual contexts in synthetic videos, testing whether MLLMs can recognize visually implausible phenomena. e.g. In the video (Sora), the watermelon breaks in the middle of the video. Is it intact or broken at the end? (Answer: It’s intact.) (Figure 1)
  • Critical Thinking Questions: Open-ended questions that ask whether MLLMs can identify abnormalities in synthetic videos, evaluating their visual reasoning. e.g. What is unusual in this video (Sora)? (Answer: The watermelon explodes, then reassembles.) (Figure 1)
while the latter two types of questions must be answered with video inputs, so that we denote them as video-dependent questions.
Model Common Sense-only Video-dependent Overall
Counterintuitive Critical Thinking
GPT-4o 100.0 46.8 15.0 45.5
InternVL3-14B 100.0 48.2 10.0 46.7
Gemini-2.5-Pro 100.0 50.2 23.3 49.8
Video-R1 100.0 52.3 16.7 50.8
Qwen2.5-VL-7B 100.0 53.1 10.0 51.0
Qwen2.5-VL-32B 100.0 52.5 13.3 51.4
Table 6: Common Sense and Video-dependent QA over VideoHallu. We divide VideoHallu into multiple categories over the question types: (a) Common Sense-only Questions, answerable via language priors without video inputs; (b) Counterintuitive Questions, probing MLLMs’ abilities in detecting counterintuitive phenomena; and (c) Critical Thinking Questions, assessing MLLMs’ ability to detect abnormalities in synthetic videos.
In Table 6, we show the evaluation breakdown by question type for six SOTA MLLMs. All models reach 100.0% accuracy on commonsense-only questions, indicating strong grounding in pre-trained knowledge. However, performance drops on counterintuitive questions (all below 55%) and further on critical thinking questions, where no model exceeds 25% accuracy, revealing major limitations in detecting and reasoning about abnormalities based on physics and commonsense.
Gemini-2.5-Pro performs best on critical thinking (23.3%), followed by Video-R1 (16.7%), suggesting some benefit from CoT prompting. However, CoT remains unreliable under language prior bias and does not consistently improve abnormality detection. Enhancing MLLMs’ critical thinking for such tasks remains an open challenge.
Counterintuitive questions typically include contextual hints, helping models locate anomalies. In contrast, critical thinking questions are open-ended, requiring models to identify and reason about abnormalities unaided, making them more vulnerable to hallucinations when their video understanding is incomplete.
Appendix D Hallucination Showcases